# Fractional Thouless pumping of solitons: a unique manifestation of bulk-edge correspondence of nonlinear eigenvalue problems

### v2

@@ III.1 Nonlinear eigenvalue of the nonlinear Hamiltonian
Figure 5: Eigenvalue’s nonlinearity of nonlinear extended RM Hamiltonian and nonlinear Thouless pumping of soliton. (a1)-(d1): The auxiliary $\lambda$-spectrum for the nonlinear extended RM Hamiltonian (1). (a2)-(d2): The anticipated position of the nonlinear excitation of soliton as a function of time over one period. The parameters are fixed at $J=1$, $\delta=0.5$, $\Delta=1$, $t_{a}=0.4$, $t_{b}=-0.1$, $T=2000\pi$, and $\omega=10^{-3}$. Specifically, (a1)-(d1) and their corresponding (a2)-(d2) panels represent interaction strengths of $g=0$, $g=1$, $g=5$ and $g=10$, respectively.
Figure 6: Effects of parameter $t_{a}$ and $t_{b}$ on the auxiliary $\lambda$-spectrum defined in Eq. (14) for the nonlinear extended RM Hamiltonian (1).The parameters are fixed at $J=1$, $\delta=0.5$, $\Delta=1$, $g=1$, $T=2000\pi$, and $\omega=10^{-3}$. Other parameters are given as (a-b): $t_{a}=0.4$, $t_{b}=-0.1$; (c-d): $t_{a}=0.5$, $t_{b}=-0.5$. (e): Effects of parameter $t_{a}$ and $t_{b}$ on the anticipated position of the nonlinear excitation of soliton as a function of time over one period.
@@ III.2 Anomalous eigenvalue’s nonlinearity of nonlinear Hamiltonian
$H_{\text{ERM}}\left(\omega,t\right)\varPsi=\omega S\left(\omega,t\right)\varPsi.$ (7)
Noted that Eq. (7) and the accompanying matrix S serve to embed auxiliary eigenvalues into the problem; As a consequence, the topological edge modes of these auxiliary eigenstates are inherited by the physical system, thereby highlighting a subtle interplay between eigenvalue nonlinearity and topology. The overlap matrix S in Eq. (7), whose structure varies with the nonlinear parameter $\omega$, is assembled as follows:
$S\left(\omega\right)=\left(\begin{array}[]{cccc}S_{0}&0&0&0\\ 0&S_{0}&0&0\\ 0&0&\ddots&0\\ 0&0&0&S_{0}\end{array}\right),$ (12)
with the diagonal block $S_{0}$ takes the form $S_{0}=1-\sigma_{z}M_{S}$ and $\sigma_{z}$ the $z$-component of Pauli matrix. Here, the scalar function $M_{S}\left(\omega\right)=M_{1}\tanh\left(\omega t\right)/\omega$ governs its dependence on the nonlinear parameter $\omega$.
@@
$P\left(\omega,t\right)=H_{\text{ERM}}\left(\omega,t\right)-\omega S\left(\omega,t\right).$ (13)
@@
Figure 7: Effects of parameter $t_{a}$ and $t_{b}$ on the auxiliary $\lambda$-spectrum defined in Eq. (14) for the nonlinear extended RM Hamiltonian (1).The parameters are fixed at $J=1$, $\delta=0.5$, $\Delta=1$, $g=7.5$, $T=2000\pi$, and $\omega=10^{-3}$. Other parameters are given as (a-b): $t_{a}=0.4$, $t_{b}=-0.1$; (c-d): $t_{a}=0.5$, $t_{b}=-0.5$. (e): Effects of parameter $t_{a}$ and $t_{b}$ on the anticipated position of the nonlinear excitation of soliton as a function of time over one period.
@@ IV Nonlinear bulk-edge correspondence
Figure 8: Band structure $\omega$-$t$ of the auxiliary eigenvalue by solving Eq. (14) at $\lambda=0$. The parameters are fixed at $g=2$, $J=1$, $\delta=0.5$, $\Delta=1$, $T=2000\pi$, and $\omega=10^{-3}$. Parameters $t_{a}$ and $t_{b}$ are given as (a): $t_{a}=0$, $t_{b}=0$; (b): $t_{a}=0.4$, $t_{b}=0.4$; (c): $t_{a}=0.4$, $t_{b}=-0.1$; (d): $t_{a}=0.4$, $t_{b}=-0.4$.
First, we analyze how the NNN hopping parameters $t_{a}$ and $t_{b}$ affect nonlinearity in the anomalous eigenvalue problem for the extended Rice-Mele Hamiltonian (Eq. 1). We specifically solve Eq. (14) at $\lambda=0$ to examine the relationship $t$$\omega$. This is achieved by computing $t$$\omega$ diagrams for varying $t_{a}$ and $t_{b}$ at fixed $\lambda=0$. Our analysis shows two key features: (i) Soliton states emerge irrespective of the specific values of $t_{a}$ and $t_{b}$, and (ii) increasing the interaction strength $g$ lowers the ground-state soliton energy. To clearly resolve two distinct soliton states, we focus our analysis on $g=2$.
@@ V Conclusion and outlook
This study primarily investigates the behavior of the nonlinear extended RM model under both linear and nonlinear eigenvalue conditions. The research explicitly demonstrates the energy spectra and phase diagrams of the system for these two distinct eigenvalue scenarios, emphasizing that the parameters $t_{a}$ and $t_{b}$ of the NNN terms can influence the system’s soliton behavior and pumping transport. Notably, we find that assigning certain specific values to the parameters $t_{a}$ and $t_{b}$ induces the emergence of fractional topological phases in the system. Moreover, when $t_{a}=-t_{b}$, the range of nonlinearity strength for adiabatic breakdown is expanded, meaning that the value of nonlinearity strength at which adiabatic breakdown begins is reduced, while the value at which it ceases is increased. Moreover, we have uncovered a fascinating phenomenon where the nonlinear parameter g can modulate the system’s topological invariants.
@@
Acknowledgements – We thank Ying Hu, Yapeng Zhang, Xuzhen Cao, Shujie Cheng, and Biao Wu for stimulating discussions and useful help. We sincerely acknowledge Yucan Yan for his dedicated efforts during the critical inception phase of this project. This work was supported by the Zhejiang Provincial Natural Science Foundation (Grant No. LZ25A040004) and the National Natural Science Foundation of China (Grants No. 12074344).

### v3

@@ III.1 Nonlinear eigenvalue of the nonlinear Hamiltonian
We remark that the fractional pumping phenomenon reported in Fig. 4(e) stems from the introduction of NNN coupling terms and is elucidated within the auxiliary eigenvalue framework in Eq. (4), rather than the conventional multi-band Wannier tracking mechanism driven by strong interactions in Ref. [55]. As explicitly demonstrated in Fig. 3(e) and Fig. 4(e), we have carefully verified that merely adjusting the nonlinearity—without including NNN coupling terms ($t_{a}=t_{b}=0$)—fails to induce fractional Thouless pumping. This conclusively shows that the predicted fractional pumping cannot be accounted for by the physical picture of solitons coupling to multi-band Wannier functions via strong interactions. Moreover, the authors in Ref. [86] report the observation of fractional Thouless pumping of solitons in a nonlinear off-diagonal Aubry–André–Harper model. Notably, this phenomenon in Ref. [86] arises even though all energy bands of the corresponding linear Hamiltonian are topologically trivial—evidence that nonlinearity can indeed induce fractional Thouless pumping of solitons. In sharp contrast, the linear Hamiltonian associated with the nonlinear eigenvalue problem $H_{\text{ERM}}\Psi=\omega S(\omega)\Psi$ is topologically nontrivial, characterized by a fractional Chern number—for instance, $\mathcal{C}=-\frac{1}{2}$ at $t_{a}=0.5$, $t_{b}=-0.5$ (red dot, Fig. 1(c)).
Figure 5: Eigenvalue’s nonlinearity of nonlinear extended RM Hamiltonian and nonlinear Thouless pumping of soliton. (a1)-(d1): The auxiliary $\lambda$-spectrum for the nonlinear extended RM Hamiltonian (1). (a2)-(d2): The anticipated position of the nonlinear excitation of soliton as a function of time over one period. The parameters are fixed at $J=1$, $\delta=0.5$, $\Delta=1$, $t_{a}=0.4$, $t_{b}=-0.1$, $T=2000\pi$, and $\omega=\omega_{\text{d}}=10^{-3}$. Specifically, (a1)-(d1) and their corresponding (a2)-(d2) panels represent interaction strengths of $g=0$, $g=1$, $g=5$ and $g=10$, respectively.
Figure 6: Effects of parameter $t_{a}$ and $t_{b}$ on the auxiliary $\lambda$-spectrum defined in Eq. (14) for the nonlinear extended RM Hamiltonian (1).The parameters are fixed at $J=1$, $\delta=0.5$, $\Delta=1$, $g=1$, $T=2000\pi$, and $\omega=\omega_{\text{d}}=10^{-3}$. Other parameters are given as (a-b): $t_{a}=0.4$, $t_{b}=-0.1$; (c-d): $t_{a}=0.5$, $t_{b}=-0.5$. (e): Effects of parameter $t_{a}$ and $t_{b}$ on the anticipated position of the nonlinear excitation of soliton as a function of time over one period.
@@ III.2 Anomalous eigenvalue’s nonlinearity of nonlinear Hamiltonian
$H_{\text{ERM}}\left(\omega_{\text{d}},t\right)\varPsi=\omega S\left(\omega\right)\varPsi.$ (7)
Noted that Eq. (7) and the accompanying matrix $S(\omega)$ serve to embed auxiliary eigenvalues into the problem; As a consequence, the topological edge modes of these auxiliary eigenstates are inherited by the physical system, thereby highlighting a subtle interplay between eigenvalue nonlinearity and topology. The overlap matrix $S(\omega)$ in Eq. (7), whose structure varies with the nonlinear parameter $\omega$, is assembled as follows:
$S\left(\omega\right)=\left(\begin{array}[]{cccc}A_{0}&0&0&0\\ 0&A_{0}&0&0\\ 0&0&\ddots&0\\ 0&0&0&A_{0}\end{array}\right),$ (12)
with the diagonal block $A_{0}$ takes the form $A_{0}=1-\sigma_{z}M_{S}$ and $\sigma_{z}$ the $z$-component of Pauli matrix. Here, the scalar function $M_{S}\left(\omega\right)=M_{1}\tanh\left(\omega t\right)/\omega$ governs its dependence on the nonlinear parameter $\omega$.
@@
$P\left(\omega,t\right)=H_{\text{ERM}}\left(\omega_{\text{d}},t\right)-\omega S\left(\omega\right).$ (13)
@@
Figure 7: Effects of parameter $t_{a}$ and $t_{b}$ on the auxiliary $\lambda$-spectrum defined in Eq. (14) for the nonlinear extended RM Hamiltonian (1).The parameters are fixed at $J=1$, $\delta=0.5$, $\Delta=1$, $g=7.5$, $T=2000\pi$, and $\omega=\omega_{\text{d}}=10^{-3}$. Other parameters are given as (a-b): $t_{a}=0.4$, $t_{b}=-0.1$; (c-d): $t_{a}=0.5$, $t_{b}=-0.5$. (e): Effects of parameter $t_{a}$ and $t_{b}$ on the anticipated position of the nonlinear excitation of soliton as a function of time over one period.
@@
We remark that we use both dynamically evolved soliton method and instantaneous soliton solution method to calculate the expectation value $\langle X\rangle=\sum_{j}j|\Psi_{j}|^{2}$ as shown in Figs. 2-7. In the first approach, we obtain the soliton at time $t$ by directly applying a given initial soliton and solving the equation for the time evolution of the wavefunction using the fourth-order Runge-Kutta method. In the second method, we determine the soliton at time $t$ through an iterative process, aiming to reach the steady state. Under adiabatic conditions, two methods converge as it is expected, providing mutual verification. Under non-adiabatic conditions related to self-intersecting bands (see Figs. 2(c1), 4(c), 5(c1) and 7(c1)) induced by the interaction, which result in discrepancies between instantaneous and dynamical soliton simulations elaborately (see Figs. 2(c2), 4(d), 5(c2) and 7(c2)), the first approach reflects physical reality in experimental implementations, while the second approach serves as a diagnostic tool for quantifying adiabaticity breakdown.
@@ IV Nonlinear bulk-edge correspondence
Figure 8: Band structure $\omega$-$t$ of the auxiliary eigenvalue by solving Eq. (14) at $\lambda=0$. The parameters are fixed at $g=2$, $J=1$, $\delta=0.5$, $\Delta=1$, $T=2000\pi$, and $\omega_{\text{d}}=10^{-3}$. Parameters $t_{a}$ and $t_{b}$ are given as (a): $t_{a}=0$, $t_{b}=0$; (b): $t_{a}=0.4$, $t_{b}=0.4$; (c): $t_{a}=0.4$, $t_{b}=-0.1$; (d): $t_{a}=0.4$, $t_{b}=-0.4$.
[…]
@@ V Conclusion and outlook
This study focuses on investigating the behavior of the nonlinear extended RM model under both linear and nonlinear eigenvalue conditions. We systematically examine the energy spectra and phase diagrams of the system under these two distinct eigenvalue scenarios, with a particular emphasis on how the parameters $t_{a}$ and $t_{b}$ of the NNN terms influence the soliton dynamics and pumping transport. A key contribution of this work lies in transforming the problem of eigenfunction nonlinearity in Eqs. (5) and (6) into a problem of eigenvalue nonlinearity in Eq. (4). Notably, we find that assigning specific values to $t_{a}$ and $t_{b}$ gives rise to the emergence of fractional topological phases in the system—an observable physical phenomenon absent in conventional frameworks ($\hat{H}\Psi=E\Psi$) and unique to the nonlinear eigenvalue problem described by Eq. (4). Furthermore, when t $t_{a}$ and $t_{b}$ , the range of nonlinearity strength enabling adiabatic breakdown broadens: the threshold for the onset of adiabatic breakdown decreases, whereas the threshold for its cessation increases. Additionally, we uncover a notable phenomenon: the nonlinear parameter $g$ modulates the system’s topological invariants.
@@
Acknowledgements – We thank Ying Hu, Yapeng Zhang, Xuzhen Cao, Shujie Cheng, and Biao Wu for stimulating discussions and useful help. We sincerely acknowledge Yucan Yan for his dedicated efforts during the critical inception phase of this project. This work was supported by the Zhejiang Provincial Natural Science Foundation (Grant No. LZ25A040004) and the National Natural Science Foundation of China (Grants No. 12574301).
Appendix A Two methods for pumping path analysis
In this Appendix A, we will provide a comprehensive description of the technical details pertaining to the computation of ground-state soliton pumping expectation values, encompassing both the instantaneous soliton and dynamical evolution methods [73, 76]. Specifically, this will include: (1) Numerical algorithms and parameter configurations for instantaneous soliton and dynamical evolution methods; (2) Annotated citations of key references supporting the methodological framework.
In the first approach, designated as the method of dynamically evolved solitons, we implement a fourth-order Runge-Kutta algorithm to numerically propagate the initial soliton profile through temporal evolution governed by the nonlinear Schrödinger equations specified in Eqs. (5) and (6). This methodology faithfully reproduces the dynamical evolution of the wavefunction under physically realistic time-dependent driving conditions, thereby providing a direct computational realization of the quantum transport process. The numerical scheme preserves the symplectic structure of the underlying physical system while maintaining a balance between computational accuracy and efficiency, with the truncation error controlled at $O(\Delta t^{4})$ through the fourth-order integration framework.
In the second approach called by the method of instantaneous solitons, an iterative self-consistent scheme is employed to converge to the steady-state solution at each time slice $t$, bypassing explicit time propagation [76]. For a given nonlinear $\omega$-dependent Hamiltonian $H(\omega_{\text{d}},t)$, at a specific time t, the iterative procedure from state $|\Psi_{n}(t)\rangle$ to $|\Psi_{n+1}(t)\rangle$ proceeds as follows:
  • First, construct the nonlinear Hamiltonian $H_{n}$ based on the given state $|\Psi_{n}(t)\rangle$;
  • Then, solve for the eigenstates $|\psi_{i}\rangleof$ $H_{n}$ with $i=1,2\dots 2N$;
  • Finally, select the eigenstate with the maximum overlap with the previous state as the new state: $|\Psi_{n+1}(t)\rangle=|\psi_{i0}\rangle$ where $\langle\Psi_{n}(t)|\psi_{i0}\rangle\geq\langle\Psi_{n}(t)|\psi_{i}\rangle$ for all i.
The iteration stops when the difference between the eigenvalues of the new state and the previous state is less than $\epsilon$ (in this work, $\epsilon=10^{-12}$ is chosen). It is crucial to note that the choice of the initial state significantly affects the final steady-state solution, hence the selection of the initial state requires extreme caution.
In the adiabatic regime ($\omega_{\text{d}}\rightarrow 0$), the two approaches converge to the same soliton trajectory, confirming numerical consistency. For nonadiabatic driving ($\omega_{\text{d}}\gg 0$), the first method faithfully reproduces experimentally observable dynamics, while the second method serves as a diagnostic tool to quantify deviations from adiabaticity by comparing instantaneous eigenstates with dynamically evolved states. We remark that, in the second approach, the wave function evolves through iterative steps as an iterative term, thereby ensuring the successful derivation of the band structure for the nonlinear wave function.